a metallurgist needs to make 24.8 oz of an alloy that is 50% gold he is going yo melt and combine one alloy that is 60% with another alloy that is 40% gold how much of each alloy should he use?
Answer by mananth(16946)

Gold
percent ---------------- quantity
Alloy I 60.00% ---------------- x Oz
Alloy II 40.00% ------ 24.8 - x Oz
Mixture 50.00% ---------------- 24.8
24.8
60.00% x + 40.00% ( 24.8 - x ) = 50.00% * 24.8
60 x + 40 ( 24.8 - x ) = 1240
60 x + 992 - 40 x = 1240
60 x - 40 x = 1240 - -992
20 x = 248
/ 20
x = 12.4 Oz 60.00% Alloy I
12.4 Oz 50.00% Alloy II

m.ananth@hotmail.ca